i am supposed to get the formula y-y1=m(x-x1) into the formula y=mx+b. the problem is: y-2=7(x-1). i have an idea how to do this, but i'm still lost.
Answer by stanbon(75887)

i am supposed to get the formula y-y1=m(x-x1) into the formula y=mx+b. the problem is:
y-2=7(x-1)
y = 7x - 7 + 2
y = 7x -5
=================
Cheers,
Stan H.